OpenStudy (anonymous):

Let f(x) =25/x The slope of the tangent line to the graph of f(x) at the point (-6,-25/6)

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

The slope of the tangent line to the graph of f(x) at the point (-6,-25/6) is ______?? The equation of the tangent line to the graph of f(x) at (-6,-25/6) is y=mx+b for m=_____?? and b=_____??

OpenStudy (anonymous):

....Let f(x) =25/x

OpenStudy (anonymous):

for (1) given that y prime is equal to the slope, find the derivative of the function,.  the result is the slope. .